# A Comprehensive Graph Pooling Benchmark: Effectiveness, Robustness and Generalizability

### v3

@@ Abstract
Graph pooling has gained attention for its ability to obtain effective node and graph representations for various downstream tasks. Despite the recent surge in graph pooling approaches, there is a lack of standardized experimental settings and fair benchmarks to evaluate their performance. To address this issue, we have constructed a comprehensive benchmark that includes 17 graph pooling methods and 28 different graph datasets. This benchmark systematically assesses the performance of graph pooling methods in three dimensions, i.e., effectiveness, robustness, and generalizability. We first evaluate the performance of these graph pooling approaches across different tasks including graph classification, graph regression and node classification. Then, we investigate their performance under potential noise attacks and out-of-distribution shifts in real-world scenarios. We also involve detailed efficiency analysis, backbone analysis, parameter analysis and visualization to provide more evidence. Extensive experiments validate the strong capability and applicability of graph pooling approaches in various scenarios, which can provide valuable insights and guidance for deep geometric learning research. The source code of our benchmark is available at https://github.com/goose315/Graph_Pooling_Benchmark.
@@ 1 Introduction
In literature, existing graph pooling approaches can be roughly divided into two categories (Bianchi & Lachi, 2024; Liu et al., 2022a), i.e., node dropping pooling (Knyazev et al., 2019; Lee et al., 2019; Ranjan et al., 2020; Ma et al., 2020b; Zhang et al., 2020a; Zhou et al., 2022; Pang et al., 2021; Bacciu et al., 2023; Zhang et al., 2020a; 2019; Song et al., 2024) and node clustering pooling approaches (Ying et al., 2018a; Bianchi et al., 2020; Duval & Malliaros, 2022; Wu et al., 2022a; Hansen & Bianchi, 2023; Tsitsulin et al., 2023a; Bianchi, 2022), based on the strategies used to simplify the graph. node dropping pooling utilizes a learnable scoring function to remove nodes with relatively low significance scores, resulting in lower computational costs, while node clustering pooling approaches typically treat graph pooling as a node clustering problem, where clusters are considered as new nodes for the coarsened graph (Liu et al., 2022b; Bianchi & Lachi, 2024). Even though graph pooling research is becoming increasingly popular, there is still no standardized benchmark that allows for an impartial and consistent comparison of various graph pooling methods. Furthermore, due to the diversity and complexity of graph datasets, numerous experimental settings have been used in previous studies, such as varied proportions of training data and train/validation/test splits (Bian et al., 2020; Hansen & Bianchi, 2023; Dwivedi et al., 2023; Xu et al., 2024b). As a result, a comprehensive and publicly available benchmark of graph pooling approaches is highly expected that can facilitate the evaluation and comparison of different approaches, ensuring the reproducibility of results and further advancing the area of graph machine learning.
@@
Open-source Material. We have made our benchmark of all these graph pooling approaches available and reproducible, and we believe our benchmark can benefit researchers in both graph machine learning and interdisciplinary fields.
@@ 2 Preliminaries
where . the process has two main principal components, i.e., reduction, and connection (Grattarola et al., 2022b). In particular, the reduction component aims to generate pooled nodes and their attributes in during the connection. The connection component computes the edges among the nodes.
@@
where denotes the node feature matrix at layer , denotes the weight matrix at the corresponding layer, is the adjacency matrix plus the identity matrix , is the degree matrix of , and is a nonlinear activation function (Kipf & Welling, 2016a). The pooling layers can be formulated as:
@@
Node Classification (Kipf & Welling, 2016a; Veličković et al., 2018; Perozzi et al., 2014). The aim of node classification is to assign semantic labels to nodes in a graph according to their attributes and relationships with different nodes. Each dataset involves a graph , consisting of nodes and their corresponding labels . is divided into a labeled set and a unlabeled set . We are required to train a graph neural network model that can predict the missing labels of nodes in using the attributes of other nodes. U-Net framework (Ronneberger et al., 2015) is widely used to incorporate pooling operations for node classification. In the encoder part, U-Net progressively applies pooling and graph convolution to downsample the graphs and extract multi-scale features. The decoder part of U-Net utilizes upsampling and graph convolution to gradually upsample the low-resolution feature maps back to the original graph size. Residual connections are employed to directly transfer the feature maps from the encoder to the decoder, facilitating the preservation of fine-grained semantics during upsampling (Ronneberger et al., 2015; Ibtehaz & Rahman, 2020; Leng et al., 2018). For more information about the U-Net, please refer to Appendix B.
@@ 3.1 Graph Pooling Approaches
Our benchmark contains 17 state-of-the-art graph pooling approaches (detailed in Table 1): TopKPool (Knyazev et al., 2019), SAGPool (Lee et al., 2019), ASAPool (Ranjan et al., 2020), PANPool (Ma et al., 2020b), COPool (Zhou et al., 2022), CGIPool (Pang et al., 2021), KMISPool (Bacciu et al., 2023), GSAPool (Zhang et al., 2020a), HGPSLPool (Zhang et al., 2019), AsymCheegerCutPool (Hansen & Bianchi, 2023), DiffPool (Ying et al., 2018a), MincutPool (Bianchi et al., 2020), DMoNPool (Tsitsulin et al., 2023a), HoscPool (Duval & Malliaros, 2022), JustBalancePool (Bianchi, 2022), SEPool (Wu et al., 2022a), and ParsPool (Song et al., 2024). More information related to the selected pooling methods can be found in the Appendix C.
@@ 3.2 Datasets
To systematically evaluate graph pooling methods, we integrate 28 datasets from different domains. For graph classification, we select eleven publicly available datasets from TUDataset (Morris et al., 2020), including seven molecules datasets, i.e., MUTAG (Debnath et al., 1991), NCI1 (Wale et al., 2008), NCI109 (Wale et al., 2008), COX2 Sutherland et al. (2003), AIDS Riesen & Bunke (2008), FRANKENSTEIN Orsini et al. (2015), and Mutagenicity Debnath et al. (1991), four bioinformatics datasets, i.e. D&D (Shervashidze et al., 2011), PROTEINS (Borgwardt et al., 2005), PROTEINS_FULL (Borgwardt et al., 2005), and ENZYMES (Schomburg et al., 2004), three social network dataset, i.e., IMDB-BINARY (IMDB-B) (Cai & Wang, 2018), IMDB-MULTI (IMDB-M) (Cai & Wang, 2018), and COLLAB (Cai & Wang, 2018). We also include a large-scale graph classification dataset, Ogbg-molpcba, from the Open Graph Benchmark (OGB) (Hu et al., 2020b). For graph regression, we choose six datasets from MoleculeNet (Wu et al., 2018) including QM7, QM8, BACE, ESOL, FreeSolv, and Lipophilicity. For node classification, we utilize three citation networks, i.e., Cora, Citeseer, and Pubmed (Yang et al., 2016), three website networks, i.e., Cornell, Texas, and Wisconsin (Pei et al., 2020), and the GitHub dataset (Rozemberczki et al., 2021). We also obtain a large-scale dataset, Ogbn-arxiv, from OGB (Hu et al., 2020b). More information about the summary statistics and description of the datasets are detailed in the Appendix D.
@@ 3.3 Evaluation Protocols
Our benchmark evaluation encompasses three key aspects, i.e., effectiveness, robustness, and generalizability. We perform a hyperparameter search for all pooling methods; detailed information can be found in Appendix E. Firstly, we conduct a performance comparison of graph pooling approaches across three tasks including graph classification, graph regression, and node classification. For graph and node classification tasks, we employ average precision for Ogbg-molpcba, and accuracy for remaining datasets as the evaluation metric. For graph regression, we use root mean square error (RMSE) for ESOL, FreeSolv, and Lipophilicity (Wu et al., 2018). Following previous research (Xu et al., 2024b), we use the area under the receiver operating characteristic (AUROC) curve to evaluate BACE, and mean absolute error (MAE) for QM7 and QM8. Secondly, our benchmark evaluates the robustness of graph pooling approaches in both graph-level and node-level tasks across two perspectives: structural robustness and feature robustness (Li & Wang, 2018). In particular, we add and drop edges of graphs to study structural robustness and mask node features to investigate feature robustness. Thirdly, we employ size-based and density-based distribution shifts to evaluate the generalizability of different pooling methods in graph-level tasks under real-world scenarios (Gui et al., 2022). We also use degree-based and closeness-based distribution shifts to assess the generalizability of different pooling methods in node-level tasks. In addition to these three views, we conduct a further analysis of these graph pooling approaches including the comparison of efficiency, visualization, and different backbone parameter choices.
@@ 4.1 Experimental Settings
All graph pooling methods in our benchmark are implemented by PyTorch (Paszke et al., 2019). Graph convolutional networks serve as the default encoders for all algorithms. The experimental setup includes a Linux server equipped with NVIDIA A100 GPUs, with an Intel Xeon Gold 6354 CPU. The software stack comprises PyTorch 1.11.0, PyTorch-geometric 2.1.0 (Fey & Lenssen, 2019), and Python 3.9.16. More details about experimental settings can be found in Appendix E.
@@ 4.2 Effectiveness Analysis
Performance on Graph Classification. To begin, we investigate the performance of different graph pooling approaches on graph classification. The results of compared approaches on seven popular datasets are recorded in Table 2. Firstly, in general, ParsPool, AsymCheegerCutPool, and COPool are the three best-performing pooling models, and the performance of all 17 pooling methods varies significantly across different datasets. No single pooling method consistently outperforms the others across all datasets. Secondly, it is noteworthy that SEPool achieves significant advantages on D&D, IMDB-B, and IMDB-M. This is because SEPool’s coding tree structure, and these datasets are characterized by high clustering coefficients Watts & Strogatz (1998), which benefits the coding tree method because the hierarchical nature of the tree can better capture and represent these localized, highly connected substructures (Wu et al., 2022a). However, SEPool also implies greater computational resource overhead, which presents challenges when processing large-scale graph data such as Ogbg-molpcba. Thirdly, the methods with the highest average accuracy are ParsPool and AsymCheegerCutPool. ParsPool can capture a personalized pooling structure for each individual graph, while AsymCheegerCutPool calculates cluster assignments based on a tighter relaxation in terms of Graph Total Variation (GTV) (Song et al., 2024; Hansen & Bianchi, 2023). These methods are flexible, and the datasets differ significantly in diameter, degree, and clustering coefficients.
@@
Performance on Node Classification. Table 4 presents the performance of various pooling methods in node classification tasks. We observe the following: Firstly, KMISPool and ParsPool demonstrate the best overall performance, significantly outperforming other methods on small-scale datasets such as Cornell, Texas, and Wisconsin. Secondly, node classification models without pooling layers achieve comparable results to most pooling methods across the majority of datasets. A potential reason for this is that pooling operations tend to lose substantial node information, which consequently weakens performance in node classification tasks. Thirdly, the scalability of ASAPool, PANPool, HGPSLPool, SEPool, and ParsPool still requires improvement, as they face memory/runtime bottlenecks, making it cannot complete training on larger datasets such as Ogbn-arxiv or GitHub.

### v4

@@ Abstract
Graph pooling has gained attention for its ability to obtain effective node and graph representations for various downstream tasks. Despite the recent surge in graph pooling approaches, there is a lack of standardized experimental settings and fair benchmarks to evaluate their performance. To address this issue, we have constructed a comprehensive benchmark that includes 17 graph pooling methods and 28 different graph datasets. This benchmark systematically assesses the performance of graph pooling methods in three dimensions, i.e., effectiveness, robustness, and generalizability. We first evaluate the performance of these graph pooling approaches across different tasks including graph classification, graph regression and node classification. Then, we investigate their performance under potential noise attacks and out-of-distribution shifts in real-world scenarios. We also involve detailed efficiency analysis, backbone analysis, parameter analysis and visualization to provide more evidence. Extensive experiments validate the strong capability and applicability of graph pooling approaches in various scenarios, which can provide valuable insights and guidance for deep geometric learning research. The source code of our benchmark is available at https://anonymous.4open.science/r/Graph_Pooling_Benchmark-8EDD.
@@ 1 Introduction
In literature, existing graph pooling approaches can be roughly divided into two categories (Bianchi & Lachi, 2024; Liu et al., 2022a), i.e., node dropping pooling (Knyazev et al., 2019; Lee et al., 2019; Ranjan et al., 2020; Ma et al., 2020b; Zhang et al., 2020a; Zhou et al., 2022; Pang et al., 2021; Bacciu et al., 2023; Zhang et al., 2020a; 2019; Song et al., 2024) and node clustering pooling approaches (Ying et al., 2018a; Bianchi et al., 2020; Duval & Malliaros, 2022; Wu et al., 2022a; Hansen & Bianchi, 2023; Tsitsulin et al., 2023a; Bianchi, 2022), based on the strategies used to simplify the graph. Node dropping pooling utilizes a learnable scoring function to guide the deletion of nodes with relatively low importance scores, resulting in lower computational costs, while node clustering pooling approaches typically treat graph pooling as a node clustering problem, where clusters are considered as new nodes for the coarsened graph (Liu et al., 2022b; Bianchi & Lachi, 2024). Even though graph pooling research is becoming increasingly popular, there is still no standardized benchmark that allows for an impartial and consistent comparison of various graph pooling methods. Furthermore, due to the diversity and complexity of graph datasets, numerous experimental settings have been used in previous studies, such as varied proportions of training data and train/validation/test splits (Bian et al., 2020; Hansen & Bianchi, 2023; Dwivedi et al., 2023; Xu et al., 2024b). As a result, a comprehensive and publicly available benchmark of graph pooling approaches is highly expected that can facilitate the evaluation and comparison of different approaches, ensuring the reproducibility of results and further advancing the area of graph machine learning.
@@
Open-source Material. We will make our benchmark of all these graph pooling approaches available and reproducible, and we believe our benchmark can benefit researchers in both graph machine learning and interdisciplinary fields.
@@ 2 Preliminaries
where . Current graph pooling methods can be classified into two categories, i.e., node dropping methods and node clustering methods (Bianchi & Lachi, 2024; Liu et al., 2022a). For node dropping methods, they focus on simplifying the graph by removing less important nodes and edges. It consists of three components, i.e., score generator for computing the importance score for each node, node selector for selecting the top- nodes with the highest importance scores, and coarsing operator for generating a new coarsened graph based on and Liu et al. (2022a). The process can be formulated as follows:
where is the feature matrix and is the adjacency matrix for layer , represents the significance scores, ranks the scores and returns the indices of the top- values in , and denotes the indices of the reserved nodes.
For node clustering methods, they focus on grouping nodes in the graph into several clusters, where each cluster represents a supernode. The graph is then simplified based on these supernodes. It consists of three components, i.e., the cluster assignment operator , which defines how nodes in the original graph are assigned to clusters in the pooled graph, mapping the nodes to the coarsened supernodes and coarsing operator to generate a new coarsened graph based on . The process can be formulated as follows:
where represents the learned cluster assignment matrix.
@@
where denotes the node feature matrix at layer , denotes the weight matrix at the corresponding layer, is the adjacency matrix plus the identity matrix , is the degree matrix of , and is a nonlinear activation function (Kipf & Welling, 2016). The pooling layers can be formulated as:
@@
Node Classification (Kipf & Welling, 2016; Veličković et al., 2018; Perozzi et al., 2014). The aim of node classification is to assign semantic labels to nodes in a graph according to their attributes and relationships with different nodes. Each dataset involves a graph , consisting of nodes and their corresponding labels . is divided into a labeled set and a unlabeled set . We are required to train a graph neural network model that can predict the missing labels of nodes in using the attributes of other nodes. U-Net framework (Ronneberger et al., 2015) is widely used to incorporate pooling operations for node classification (Wu et al., 2022a; Song et al., 2024; Zhang et al., 2021). Figure 2 shows the overview of the U-Net framework. In the encoder part, U-Net progressively applies pooling and graph convolution to downsample the graphs and extract multi-scale features. The decoder part of U-Net utilizes upsampling and graph convolution to gradually upsample the low-resolution feature maps back to the original graph size. Residual connections are employed to directly transfer the feature maps from the encoder to the decoder, facilitating the preservation of fine-grained semantics during upsampling (Ronneberger et al., 2015; Ibtehaz & Rahman, 2020; Leng et al., 2018). Pooling not only simplifies the graph’s complexity but also provides the model with multi-scale feature representation capabilities.
In particular, in the downsampling path, the input feature matrix is first subjected to graph convolution, where the product of the adjacency matrix and the feature matrix, along with the weighted sum of the weight matrix and the bias term, yields the activated feature matrix . Next, a pooling operation is applied, reducing the number of nodes by selecting those with higher scores, thereby transforming the original feature matrix and adjacency matrix into a smaller feature matrix and a corresponding adjacency matrix . In the upsampling path, the pooled feature matrix is first upsampled to restore the original number of nodes, generating a new feature matrix . Then, the restored feature matrix is concatenated with the corresponding feature matrix from the downsampling path, forming the merged feature matrix . Finally, the merged feature matrix undergoes another graph convolution, resulting in the output feature matrix .
@@ 3.1 Graph Pooling Approaches
Graph pooling plays a significant role in graph mining learning. First, its primary function is to compress node representations into a smaller graph or a single vector, which can reduce computational complexity. This is particularly important for processing large-scale graph data with scalability (Knyazev et al., 2019; Ju et al., 2024). Second, graph pooling generates global graph representations by aggregating local node information, enabling better capture of substructures and motifs (Song et al., 2024). Third, hierarchical pooling methods, such as node clustering pooling and node drop pooling, coarsen the graph step-by-step, extracting features at multiple levels. This helps capture complex structures and multi-level information, thereby enhancing the model’s expressive power (Zhou et al., 2022; Bacciu et al., 2023). Finally, graph pooling improves the model’s generalization ability to unseen graph data by reducing graph complexity and noise, enhancing robustness, especially in out-of-distribution (OOD) scenarios (Duval & Malliaros, 2022; Ju et al., 2024). Therefore, we introduce this benchmark for evaluating and understand graph pooling approaches.
Our benchmark contains 17 state-of-the-art graph pooling approaches (detailed in Table 1): TopKPool (Knyazev et al., 2019), SAGPool (Lee et al., 2019), ASAPool (Ranjan et al., 2020), PANPool (Ma et al., 2020b), COPool (Zhou et al., 2022), CGIPool (Pang et al., 2021), KMISPool (Bacciu et al., 2023), GSAPool (Zhang et al., 2020a), HGPSLPool (Zhang et al., 2019), AsymCheegerCutPool (Hansen & Bianchi, 2023), DiffPool (Ying et al., 2018a), MincutPool (Bianchi et al., 2020), DMoNPool (Tsitsulin et al., 2023a), HoscPool (Duval & Malliaros, 2022), JustBalancePool (Bianchi, 2022), SEPool (Wu et al., 2022a), and ParsPool (Song et al., 2024). Then, we introduce the details of these graph pooling approaches:
@@ 3.2 Datasets
To systematically evaluate graph pooling methods, we integrate 28 datasets from different domains. For graph classification, we select eleven publicly available datasets from TUDataset (Morris et al., 2020), including seven molecules datasets, i.e., MUTAG (Debnath et al., 1991), NCI1 (Wale et al., 2008), NCI109 (Wale et al., 2008), COX2 (Sutherland et al., 2003), AIDS (Riesen & Bunke, 2008), FRANKENSTEIN (Orsini et al., 2015), and Mutagenicity (Debnath et al., 1991), four bioinformatics datasets, i.e. D&D (Shervashidze et al., 2011), PROTEINS (Borgwardt et al., 2005), PROTEINS_FULL (Borgwardt et al., 2005), and ENZYMES (Schomburg et al., 2004), three social network dataset, i.e., IMDB-BINARY (IMDB-B) (Cai & Wang, 2018), IMDB-MULTI (IMDB-M) (Cai & Wang, 2018), and COLLAB (Cai & Wang, 2018). We also include a large-scale graph classification dataset, Ogbg-molpcba, from the Open Graph Benchmark (OGB) (Hu et al., 2020b). For graph regression, we choose six datasets from MoleculeNet (Wu et al., 2018) including QM7, QM8, BACE, ESOL, FreeSolv, and Lipophilicity. For node classification, we utilize three citation networks, i.e., Cora, Citeseer, and Pubmed (Yang et al., 2016), three website networks, i.e., Cornell, Texas, and Wisconsin (Pei et al., 2020), and the GitHub dataset (Rozemberczki et al., 2021). We also obtain a large-scale dataset, Ogbn-arxiv, from OGB (Hu et al., 2020b). More information about the summary statistics and description of the datasets are detailed in the Appendix A.
@@ 3.3 Evaluation Protocols
Our benchmark evaluation encompasses three key aspects, i.e., effectiveness, robustness, and generalizability. We perform a hyperparameter search for all pooling methods; detailed information can be found in Appendix B. Firstly, we conduct a performance comparison of graph pooling approaches across three tasks including graph classification, graph regression, and node classification. For graph and node classification tasks, we employ average precision for Ogbg-molpcba, and accuracy for remaining datasets as the evaluation metric. For graph regression, we use root mean square error (RMSE) for ESOL, FreeSolv, and Lipophilicity (Wu et al., 2018). Following previous research (Xu et al., 2024b), we use the area under the receiver operating characteristic (AUROC) curve to evaluate BACE, and mean absolute error (MAE) for QM7 and QM8. Secondly, our benchmark evaluates the robustness of graph pooling approaches in both graph-level and node-level tasks across two perspectives: structural robustness and feature robustness (Li & Wang, 2018). In particular, we add and drop edges of graphs to study structural robustness and mask node features to investigate feature robustness. Thirdly, we employ size-based and density-based distribution shifts to evaluate the generalizability of different pooling methods in graph-level tasks under real-world scenarios (Gui et al., 2022). We also use degree-based and closeness-based distribution shifts to assess the generalizability of different pooling methods in node-level tasks. In addition to these three views, we conduct a further analysis of these graph pooling approaches including the comparison of efficiency, visualization, and different backbone parameter choices.
@@ 4.1 Experimental Settings
All graph pooling methods in our benchmark are implemented by PyTorch (Paszke et al., 2019). Graph convolutional networks serve as the default encoders for all algorithms. The experimental setup includes a Linux server equipped with NVIDIA A100 GPUs, with an Intel Xeon Gold 6354 CPU. The software stack comprises PyTorch 1.11.0, PyTorch-geometric 2.1.0 (Fey & Lenssen, 2019), and Python 3.9.16. More details about experimental settings can be found in Appendix B.
@@ 4.2 Effectiveness Analysis
Performance on Graph Classification. To begin, we investigate the performance of different graph pooling approaches on graph classification. The results of compared approaches on seven popular datasets are recorded in Table 2. Firstly, in general, ParsPool, AsymCheegerCutPool, and COPool are the three best-performing pooling models, and the performance of all 17 pooling methods varies significantly across different datasets. No single pooling method consistently outperforms the others across all datasets. Secondly, it is noteworthy that SEPool achieves significant advantages on D&D, IMDB-B, and IMDB-M. This is because SEPool’s coding tree structure, and these datasets are characterized by high clustering coefficients (Watts & Strogatz, 1998), which benefits the coding tree method because the hierarchical nature of the tree can better capture and represent these localized, highly connected substructures (Wu et al., 2022a). However, SEPool also implies greater computational resource overhead, which presents challenges when processing large-scale graph data such as Ogbg-molpcba. Thirdly, the methods with the highest average accuracy are ParsPool and AsymCheegerCutPool. ParsPool can capture a personalized pooling structure for each individual graph, while AsymCheegerCutPool calculates cluster assignments based on a tighter relaxation in terms of Graph Total Variation (GTV) (Song et al., 2024; Hansen & Bianchi, 2023). These methods are flexible, and the datasets differ significantly in diameter, degree, and clustering coefficients. Fourthly, GSAPool demonstrates superior performance on datasets with small-scale graphs, such as Ogbg-molpcba (Avg. nodes=26.0, Avg. edges=27.5) and MUTAG (Avg. nodes=17.9, Avg. edges=39.6). GSAPool combines structural information (SBTL) and node feature information (FBTL) to generate the final pooling topology. In terms of structural information, SBTL employs GCNConv, while for feature information, FBTL directly processes node features using MLP. GCNConv accurately identifies critical structural nodes (e.g., high-degree hub nodes) through neighborhood aggregation. The limited number of nodes enhances the distinctiveness of individual node features (e.g., atomic attributes in molecular graphs), enabling MLP to efficiently extract their importance. Additionally, on small-scale graphs, structural information (e.g., node degree and adjacency relations) and feature information (e.g., node attributes) are typically more pronounced and easier to capture, leading to better performance. ParsPool excels on datasets with large-scale graphs, such as D&D (Avg. nodes=284.3, Avg. edges=715.7) and COLLAB (Avg. nodes=74.5, Avg. edges=2457.2). The algorithm combines graph topology with continuous feature values through a learnable edge score matrix, generating clusters via three operations: DOM, EXP, and GEN. DOM selects key edges to anchor initial cluster centers, EXP dynamically absorbs neighboring nodes through expansion, and GEN produces a soft assignment matrix. This process adaptively partitions clusters based on local structures and global connectivity patterns without predefined cluster numbers or pooling ratios. For large-scale graphs with dense edges (e.g., COLLAB), the learnable edge scores enable gradient-based optimization to identify high-order semantic clusters while sparsifying non-critical edges. For graphs with numerous nodes (e.g., D&D), the row-wise max-indexing in DOM and neighborhood expansion in EXP efficiently aggregate nodes with linear complexity, avoiding the performance bottlenecks of traditional global clustering methods.
@@
Performance on Node Classification. Table 4 presents the performance of various pooling methods in node classification tasks. We observe the following: Firstly, KMISPool and ParsPool demonstrate the best overall performance, significantly outperforming other methods on small-scale datasets such as Cornell, Texas, and Wisconsin. Secondly, node classification models without pooling layers achieve comparable results to most pooling methods across the majority of datasets. A potential reason for this is that pooling operations tend to lose substantial node information, which consequently weakens performance in node classification tasks. Thirdly, the scalability of ASAPool, PANPool, HGPSLPool, SEPool, and ParsPool still requires improvement, as they face memory/runtime bottlenecks, making it cannot complete training on larger datasets such as Ogbn-arxiv or GitHub. Fourthly, for datasets with strong connectivity between nodes, such as Ogbn-arxiv (Avg. degree=13.7) and Github (Avg. degree=15.3), GSAPool performs exceptionally well due to its emphasis on extracting structural information through GCNConv. Conversely, for datasets with weaker connectivity, such as PubMed (Avg. degree=4.5), Texas (Avg. degree=3.2), and Wisconsin (Avg. degree=3.7), KMISPool excels. KMISPool demonstrates superior performance on the Cora dataset, which has the highest average clustering coefficient (Avg. CC=0.24). A high clustering coefficient indicates strong community structure among nodes. The k-MIS algorithm selects nodes that are more than k hops apart as centroids, ensuring their uniform distribution across the graph and preventing over-sampling or omission in specific regions. This uniform node selection strategy is particularly crucial in graphs with weaker connectivity, where traditional pooling methods may suffer from information loss or uneven sampling due to insufficient local connections.